OpenStudy (anonymous):

show all work simplifying 1/x + 3/8 = 1/2?

jimthompson5910 (jim_thompson5910):

the denominators are x, 8 and 2 The LCD (lowest common denominator) is 8x

jimthompson5910 (jim_thompson5910):

Multiply both sides by the LCD to clear out the fractions \[\Large \frac{1}{x}+\frac{3}{8} = \frac{1}{2}\] \[\Large {\color{red}{8x*}}\left(\frac{1}{x}+\frac{3}{8}\right) = {\color{red}{8x*}}\frac{1}{2}\] \[\Large {\color{red}{8x*}}\left(\frac{1}{x}\right)+{\color{red}{8x*}}\left(\frac{3}{8}\right) = {\color{red}{8x*}}\frac{1}{2}\] \[\Large \frac{8x}{x}+\frac{24x}{8} = \frac{8x}{2}\] \[\Large 8 + 3x = 4x\] I'll let you finish solving for x.